# What are the Essential Factors in Crafting Effective Long Context Multi-Hop Instruction Datasets? Insights and Best Practices

### v1

@@ Appendix A Data Construction Details
The construction of long-text multi-hop question-and-answer datasets is based on a structured approach leveraging pre-trained document corpora. This section outlines the methodology used for data collection, processing, and validation across multiple domains and languages.
A.1 Source Data Overview
@@
The primary source of long-text data is a pre-trained document corpus that spans nine distinct domains. The corpus includes data from both Chinese and English sources, ensuring a comprehensive multilingual dataset. The domains covered are:
@@
Additionally, inspired by Kim et al. (2023) and Chen et al. (2024a), CoT has the ability to bring powerful performance improvements to the instruction tuning. What’s more, as shown in Figure 18, after adding CoT, the performance of the model has indeed improved. Therefore, in all our data synthesis processes, the answer contains a reasoning path. Furthermore, since LLMs often cannot fit all the document information that is extremely long documents, we perform truncation segmentation on the documents input to the model. After generating the sample, refill the document with other documents to a fixed length.
A.2 Multi-hop Question and Answer Data Construction
@@ A.2.1 Dataset Curation
A.2.2 Quality Verification Agent
The first module in our framework is Quality Verification Agent, which ensures that the generated questions and answers meet a certain standard of quality. We use InternLM2-20B (Cai et al., 2024) as backbone and set the quality score threshold to 8.5. Moreover, the prompt are as follows:
Suppose you are a professional annotator, and you need to annotate the generated questions, rationales and answers according to the context. Specifically, your tasks are as follows: First, determine whether the questions and answers are in documents provided in context. Then, you need to determine whether the problem is a multi-hop problem, using multi-hop logic. At the same time, you need to judge whether the question conforms to commonsense logic. Does the question conform to common sense in a normal context? Is the logic smooth? In addition, you need to rate the overall data quality from three aspects: logical rationality and fluency, question complexity, and answer clarity. All scores are between 0 and 10. Before giving an annotation, you need to give your rationale. [[DOCUMENTS]] {chunk} [[QUESTION]] {question} [[ANSWER]] {answer} Finally, you should give me an overall quality mark in the format: “‘{”in_document”: BOOL, ”domain_similarity”: NUMBER, ”quality”: NUMBER}“‘
A.2.3 Single-hop Question Generation Agent
@@
The document content is as follows: {chunk}
Extract the questions contained in the above document, and the extracted questions should meet the following conditions: No pictorial information should be included in the extracted questions; No referential information should be included in the extracted questions; Ensure the completeness of the extracted questions; if they are multiple-choice questions, provide corresponding option information, remove line breaks, and place the question body in a single question; If the document contains concepts such as numbers, time, people, or places, questions that involve this information must be extracted; The extracted questions should be presented in a parseable list format, such as [”xxx”, ”xxx”]. If there are no valuable questions, output an empty list []; Try to extract as many valuable questions as possible, but do not include duplicate questions; Extract no more than three questions; Extracted questions:
Based on the questions extracted, the prompt of answer generation is as follows:
@@
A.2.4 Multiple Question Sampling
@@
A.2.5 Multi-hop Question Merging Agent
@@
Based on the given two question answer pairs, synthesize up to one question answer pair that matches the real scenario. The synthesized question answer pair should meet the following conditions: If both questions and answers are time related, a comparative question can be synthesized to compare the order in which two events occur; If both questions and answers are related to the character, it can be synthesized to determine which character better fits the description of the composite question; The synthesized answer should provide the corresponding reasoning process, and the synthesized answer should make as much use of the content in the given two answers as possible; Do not arbitrarily change the original information of two questions and answers; The generated questions and answers are strictly output in JSON format using {”question”: xxx, ”answer”: xxx}. Synthesized question answer pair should not have any line breaks; The correct answers to two questions are as follows: {qa1} {qa2} The synthesized question answer pair is:
Figure 15: Analysis of the impact of different training dataset size on the average accuracy score.
@@
Appendix B Highest Quality Strategies Details
To achieve the highest quality data, we deliberately prioritize the use of GPT-4o as the backbone for all processes, fully disregarding cost constraints. This decision is driven by the understanding that ensuring the best data quality is paramount for the success of our project. Furthermore, to maintain and enhance performance during the exploration phase, we implement a comprehensive range of strategies aimed at maximizing the data retention rate.
@@ Appendix B Highest Quality Strategies Details
Appendix C Instruction Tuning Experiments Details
C.1 Training Details
All models were trained using 64 A800*80G GPUs with the DeepSpeed+ZeRO-1 framework. The maximum sequence length was set to 32K, with any sequences exceeding this length truncated from the right. The training process utilized the Adam optimizer with a learning rate of $3\times 10^{-5}$, $\beta_{1}=0.9$, and $\beta_{2}=0.95$.
@@ C.1 Training Details
C.2 Evaluation Details
Based on the methodology proposed by Bai et al. (2024a), evaluating Token F1 using a model optimized through Chain of Thought (CoT) (Wei et al., 2022) reasoning proves to be challenging. To address this limitation, we employ GPT-4 as a consistency evaluator. Our testing demonstrates that the error rate of GPT-4 in this role remains consistently low, with deviations falling within a 2% margin. The corresponding prompt used is outlined below:
Suppose you are a professional annotator. Given the result predicted by a model, you need to annotate whether the “[[PREDICTION]]“ is consistent with the given “[[REFERENCE]]“ based on the “[[QUESTION]]“. [[QUESTION]] {question} [[PREDICTION]] {predictions} [[REFERENCE]] {answer} Finally, you should give me an annotation in the format: “‘ { ”short_pred_answer”: ”xxx”, ”predict_consistency”: BOOL } “‘
@@ C.2 Evaluation Details
Figure 18: The case study of the effectiveness of LongMIT.
Appendix D Case study
@@
  • Impact of Instruction Quality: As illustrated in Figure 18, models trained with high-quality multi-hop instruction data, specifically the LongMIT dataset, exhibit enhanced logical reasoning capabilities. These models are better equipped to process and analyze extensive textual information, enabling them to derive more accurate and reliable reasoning. In contrast, models trained using traditional, lower-quality instruction data, such as LongAlign Bai et al. (2024a), demonstrate a reduced capacity for logical reasoning. This comparison underscores the importance of the quality of training data in developing models that can effectively handle complex reasoning tasks, especially when dealing with long and intricate texts.
@@
  • Role of Rationale Incorporation in Training: Furthermore, as depicted in Figure 17, our analysis reveals that the inclusion of additional rationales during the training process significantly enhances the model’s ability to focus on relevant information within long texts and make precise inferences. This finding is particularly evident when comparing models that underwent Chain-of-Thought (CoT) Wei et al. (2022) training with those that did not. Specifically, models that lacked CoT training tend to falter during inference, often generating erroneous outputs, such as the completely incorrect answer ”1976”. On the other hand, models that were fine-tuned with CoT training not only demonstrate a coherent logical reasoning process but also consistently arrive at the correct answer, ”1065”. This result highlights the critical role of rationale-based training in improving the model’s reasoning accuracy and its ability to tackle complex inferential challenges.

### v2

@@ C.1 Dataset Construction Pipeline Discussion
The construction of long-context multi-hop question-and-answer datasets is based on a structured approach leveraging pre-trained document corpora. This section outlines the methodology used for data collection, processing, and validation across multiple domains and languages.
C.2 Source Data Overview
@@
The primary source of long-text data is a pre-trained document corpus that spans nine distinct and most widely-used domains (Biancofiore et al., 2024). Inspired by Qin et al. (2025); Zhang et al. (2024b), the corpus includes data from the most widely-used bilingual sources (viz. Chinese and English), ensuring a comprehensive multilingual dataset. The domains covered are:
@@
Additionally, chain-of-thought (CoT) has become a widely-used effective technique for various tasks (Kim et al., 2023; Chen et al., 2024b). Inspired by this, CoT has the ability to bring powerful performance improvements to the instruction tuning. What’s more, as shown in Figure 13, after adding CoT, the performance of the model has indeed improved, which following the recent reasoning conclusion (Guo et al., 2025; Zhuang et al., 2023; Chen et al., 2024a, 2025b). Therefore, in all our data synthesis processes, the answer contains a reasoning path. Furthermore, since LLMs often cannot fit all the document information that is extremely long documents, we perform truncation segmentation on the documents input to the model. After generating the sample, refill the document with other documents to a fixed length.
C.3 Instruction Dataset Construction
To expand the domain coverage and handle longer contexts, we extend the instruction fine-tuning data across 9 domains and 2 languages. All base documents are sourced from pre-trained datasets to prevent data leakage. Our Long Multi-hop Instruction-Tuning dataset (LongMIT) results in a retention rate of over 90% in GPT-4o verification in 200 sampled samples, confirming the high quality and generalizability of our pipeline. To balance the cost and effectiveness of generating data, LongMIT are generated based on Qwen2-72B-Instruct and verified based on InternLM2-20B. We conduct a detailed statistical analysis of sample size and token consumption across various datasets.
Dataset Name Sample Size Token Size LongAlpaca (Chen et al., 2024c) 12.00k 0.11B LongAlign (Bai et al., 2024a) 9.89k 0.17B ChatQA2 (Xu et al., 2024a) 128.00k 1.22B NQ (Kwiatkowski et al., 2019) 315.20k 4.56B LongMIT (Ours) 64.40k 5.07B
Table 3: The statistics results of different datasets.
Moreover, Table 3 presents the sample and token sizes. Our token count is comparable to that of NQ, but our dataset, while containing fewer samples, outperforms NQ by over 10%. Notably, despite NQ’s larger sample size, its performance lags behind datasets with fewer samples, such as LongAlpaca and LongAlign.
C.4 Multi-hop Question and Answer Data Construction
@@ C.4.1 Dataset Curation
C.4.2 Quality Verification Agent
The first module in our framework is the Quality Verification Agent, which ensures that the generated questions and answers meet a certain standard of quality. We use InternLM2-20B (Cai et al., 2024) as the backbone and set the quality score threshold to 8.5. Moreover, the prompts are as follows:
Suppose you are a professional annotator, and you need to annotate the generated questions, rationales, and answers according to the context. Specifically, your tasks are as follows: First, determine whether the questions and answers are in documents provided in context. Then, you need to determine whether the problem is a multi-hop problem, using multi-hop logic. At the same time, you need to judge whether the question conforms to commonsense logic. Does the question conform to common sense in a normal context? Is the logic smooth? In addition, you need to rate the overall data quality from three aspects: logical rationality and fluency, question complexity, and answer clarity. All scores are between 0 and 10. Before giving an annotation, you need to give your rationale. [[DOCUMENTS]] {chunk} [[QUESTION]] {question} [[ANSWER]] {answer}
Finally, you should give me an overall quality mark in the format: “‘{"in_document": BOOL, "domain_similarity": NUMBER, "quality": NUMBER}“‘
C.4.3 Single-hop Question Generation Agent
@@
The document content is as follows: {chunk} Extract the questions contained in the above document, and the extracted questions should meet the following conditions: No pictorial information should be included in the extracted questions; No referential information should be included in the extracted questions; Ensure the completeness of the extracted questions; if they are multiple-choice questions, provide corresponding option information, remove line breaks, and place the question body in a single question; If the document contains concepts such as numbers, time, people, or places, questions that involve this information must be extracted; The extracted questions should be presented in a parseable list format, such as ["xxx", "xxx"]. If there are no valuable questions, output an empty list []; Try to extract as many valuable questions as possible, but do not include duplicate questions; Extract no more than three questions; Extracted questions:
Based on the questions extracted, the prompt for answer generation is as follows:
@@
C.4.4 Multiple Question Sampling
@@
C.4.5 Multi-hop Question Merging Agent
@@
Based on the given two question-answer pairs, synthesize up to one question-answer pair that matches the real scenario. The synthesized question-answer pair should meet the following conditions: If both questions and answers are time-related, a comparative question can be synthesized to compare the order in which two events occur; If both questions and answers are related to the character, it can be synthesized to determine which character better fits the description of the composite question; The synthesized answer should provide the corresponding reasoning process, and the synthesized answer should make as much use of the content in the given two answers as possible; Do not arbitrarily change the original information of two questions and answers; The generated questions and answers are strictly output in JSON format using {"question": xxx, "answer": xxx}. Synthesized question-answer pairs should not have any line breaks;
The correct answers to two questions are as follows: {qa1} {qa2} The synthesized question-answer pair is:
InterLM-2.5-7B-Enhance InterLM-2.5-7B-Enhance + LongMIT 4k 8k 16k 32k 128k 4k 8k 16k 32k 128k S-NIAH Subtask-1 99.00 99.00 100.00 100.00 16.00 99.00 99.00 99.00 99.00 97.00 S-NIAH Subtask-2 100.00 99.00 100.00 99.00 97.00 100.00 100.00 100.00 100.00 100.00 S-NIAH Subtask-3 99.00 98.00 99.00 99.00 100.00 99.00 99.00 99.00 99.00 100.00 MK-NIAH Subtask-1 97.00 98.00 97.00 88.00 58.00 100.00 100.00 98.00 99.00 90.00 MK-NIAH Subtask-2 99.00 99.00 96.00 81.00 28.00 99.00 100.00 100.00 95.00 63.00 MK-NIAH Subtask-3 95.00 90.00 56.00 14.00 0.00 96.00 91.00 70.00 33.00 2.00 MV-NIAH 99.25 99.50 99.50 94.50 84.50 99.00 99.00 97.00 93.50 89.50 MQ-NIAH 98.00 98.75 97.50 94.00 86.00 100.00 100.00 100.00 99.25 94.25 VT 91.20 91.80 98.60 97.40 0.00 96.60 97.80 98.80 95.60 94.20 FWE 85.33 87.00 84.67 91.00 71.67 86.00 89.00 86.67 90.67 78.33 CWE 83.40 67.10 34.70 26.40 0.10 75.60 40.40 8.50 6.10 0.20 Single Hop QA 90.00 80.00 81.00 75.00 42.00 92.00 84.00 82.00 80.00 58.00 Multi Hop QA 70.00 67.00 64.00 53.00 35.00 73.00 72.00 68.00 63.00 47.00 Average 92.78 90.32 85.23 77.87 47.56 93.48 90.09 85.15 81.01 70.27
Table 4: The evaluation performance on Ruler (Hsieh et al., 2024) benchmark based on LongMIT extended to 128K.
Appendix D Highest Quality Strategy Details
D.1 Different Generation Strategy
D.1.1 Generation Strategy Defination
LongMIT+Best-Strategy means that add all strategies that yield better performance but may incur higher costs. These include using GPT-4o as the backbone model, incorporating additional rationales, merging questions with corresponding documents, and other related techniques. More details are described in Sec. D.2.
Self-Instruct strategy involves prompting GPT-4o to autonomously generate questions and corresponding answers based on the provided document, leveraging its self-instruction capabilities for required outputs.
D.1.2 Generation Strategy Discussion
As shown in Figure 8, the quality of the data synthesized using Qwen2 within the MIMG framework significantly exceeds that generated by GPT-4 using the Self-Instruct strategy. This illustrates the effectiveness of our framework, even more effective than backbone replacement. Furthermore, as illustrated in Table 1, our approach consistently outperforms prior efforts utilizing GPT-3.5, GPT-4, and even human-created data, which demonstrates and further proves the effectiveness of our framework.
In addition, GPT4 is often impractical for training data synthesis. It is also important to note that the large-scale training data, distilled from only 4 billion tokens using GPT-4, incurs a cost exceeding \$10,000, rendering such an approach impractical for many applications.
Figure 10: Analysis of the impact of different training dataset sizes on the average accuracy score.
@@ D.1.2 Generation Strategy Discussion
D.2 Implementation Details
To achieve the highest quality data, we deliberately prioritize the use of GPT-4o as the backbone for all processes, fully disregarding cost constraints. This decision is driven by the understanding that ensuring the best data quality is paramount for the success of our project. Furthermore, to maintain and enhance performance during the exploration phase, we implement a range of strategies aimed at maximizing the data retention rate.
@@ D.2 Implementation Details
High-quality Diversity Multi-hop Self-Instruct 61.3 53.4 33.1 MIMG 94.8 88.2 94.8       w/o MQMA 85.0 85.7 45.7       w/o MQS 92.8 65.4 89.4       w/o SQGA 82.4 69.4 73.2       w/o QVA 78.3 88.1 91.0
Table 5: Results of the ablation study on MIMG, reporting the data quality score as the average of three independent manual evaluations.
Model inst_level_loose_acc inst_level_strict_acc prompt_level_loose_acc prompt_level_strict_acc Llama-3-8B-ProLong 0.1259 0.1199 0.0610 0.0536     +ChatQA2 0.0156 0.0132 0.0074 0.0037     +LongMIT 0.1583 0.1463 0.0795 0.0684
Table 6: The instruction-following capabilities on IFEval for different instruction datasets.
High-quality Diversity Multi-hop InternLM2-Chat-20B (Cai et al., 2024) Self-Instruct 51.0 44.0 52.0 MIMG 86.0 66.0 80.0 Qwen-72B-Instruct (Yang et al., 2024) Self-Instruct 61.3 53.4 33.1 MIMG 94.8 88.2 94.8 GPT-4o (Achiam et al., 2023) Self-Instruct 86.0 81.0 66.0
Table 7: The human-annotated quality score for different LLMs and strategies.
D.3 Ablation Analysis
We analyze the contributions of various agent components to the quality of human-annotated data. As summarized in Table 5, each component positively impacts the overall quality. The Multi-hop Question Merger Agent notably improves the multi-hop quality, while the Multiple Question Sampling mechanism increases data diversity. The Single-hop Question Generation Agent is crucial for enhancing both quality and diversity. Lastly, the Quality Verification Agent acts as a safeguard, ensuring a lower bound of model quality and further improving data integrity.
D.4 More Backbone Exploration
To assess the data generation quality across additional backbone architectures, we manually labeled 100 samples for both open-source and closed-source models. As reported in Table 7, integrating MIMG with InternLM2 yields a marked improvement in output quality. Moreover, LongMIT surpasses the combination of GPT-4o and Self-Instruct on our quality metrics. We will elaborate on these findings in the next version.
D.5 Instruction Capabilities
As MIMG is designed to enhance models’ ability to follow instructions over extended contexts. Accordingly, it is essential both to improve instruction-following performance and to preserve this ability in long-context settings. Because there is currently no benchmark dedicated to long-context instruction following, we evaluated instruction adherence using two established short-context benchmarks: IFEval and ArenaHard.
As shown in Table 6, remarkably, the model trained with LongMIT not only achieved substantial gains in handling long contexts but also outperformed the base model on the IFEval benchmark. In contrast, training on ChatQA2 led to a pronounced decline in instruction-following performance on IFEval.
In addition, as presented in Table 8, on the more demanding short-context instruction-following benchmark ArenaHard, our method continues to outperform ChatQA2 by a substantial margin. Moreover, the improvements to long-context processing do not compromise its effectiveness on challenging instruction-following tasks.
ArenaHard Score 95% Confidence Interval Llama-3-8B-ProLong 7.2 (1.2, 1.2)1.21.2(-1.2,\;1.2)( - 1.2 , 1.2 )       +ChatQA2 4.4 (0.9, 0.8)0.90.8(-0.9,\;0.8)( - 0.9 , 0.8 )       +LongMIT 6.7 (1.1, 1.0)1.11.0(-1.1,\;1.0)( - 1.1 , 1.0 )
Table 8: The instruction-following capabilities on ArenaHard for different instruction datasets.
Figure 12: The case study of the effectiveness of LongMIT.
Appendix E Instruction Tuning Experiments Details
E.1 Training Details
All models were trained using 64 A800*80G GPUs with the DeepSpeed+ZeRO-1 framework. The maximum sequence length was set from 4K to 128K, with any sequences exceeding this length truncated from the right. The training process utilized the Adam optimizer with a learning rate of $3\times 10^{-5}$, $\beta_{1}=0.9$, and $\beta_{2}=0.95$.
@@ E.1 Training Details
E.2 Evaluation Details
Based on the methodology proposed by Bai et al. (2024a), evaluating Token F1 using a model optimized through Chain of Thought (CoT) (Wei et al., 2022) reasoning proves to be challenging (Chen et al., 2025a). To address this limitation, we employ GPT-4 as a consistency evaluator. Our testing demonstrates that the error rate of GPT-4 in this role remains consistently low, with deviations falling within a 2% margin. The corresponding prompt used is outlined below:
Suppose you are a professional annotator. Given the result predicted by a model, you need to annotate whether the “[[PREDICTION]]“ is consistent with the given “[[REFERENCE]]“ based on the “[[QUESTION]]“. [[QUESTION]] {question} [[PREDICTION]] {predictions} [[REFERENCE]] {answer} Finally, you should give me an annotation in the format: “‘ { "short_pred_answer": "xxx", "predict_consistency": BOOL } “‘
@@ E.2 Evaluation Details
Appendix F Case study
@@
  • Impact of Instruction Quality: As illustrated in Figure 13, models trained with high-quality multi-hop instruction data, specifically the LongMIT dataset, exhibit enhanced logical reasoning capabilities. These models are better equipped to process and analyze extensive textual information, enabling them to derive more accurate and reliable reasoning. In contrast, models trained using traditional, lower-quality instruction data, such as LongAlign (Bai et al., 2024a), demonstrate a reduced capacity for logical reasoning. This comparison underscores the importance of the quality of training data in developing models that can effectively handle complex reasoning tasks, especially when dealing with long and intricate texts.
@@
  • Role of Rationale Incorporation in Training: Furthermore, as depicted in Figure 12, our analysis reveals that the inclusion of additional rationales during the training process significantly enhances the model’s ability to focus on relevant information within long texts and make precise inferences. This finding is particularly evident when comparing models that underwent Chain-of-Thought (CoT) (Wei et al., 2022) training with those that did not. Specifically, models that lacked CoT training tend to falter during inference, often generating erroneous outputs, such as the completely incorrect answer "1976". On the other hand, models that were fine-tuned with CoT training not only demonstrate a coherent logical reasoning process but also consistently arrive at the correct answer, "1065". This result highlights the critical role of rationale-based training in improving the model’s reasoning accuracy and its ability to tackle complex inferential challenges.
Appendix G Discussion about Long Code Data Generation
In both coding and mathematical settings, it is essential to perform multi-step reasoning across source files and scholarly articles. To investigate this requirement, we evaluate our framework on a set of coding tasks. The resulting code–question–answer dataset exhibited a high level of quality, with approximately 80% of entries rated as satisfactory. Furthermore, by consolidating related questions, we are capable to prompt the advanced LLM (GPT-4o) to generate code-generation queries of competition caliber.
Below, we illustrate a simple example in which the context consists of code documents from the “torch.nn” module. First, the Single-Hop Question Generation Agent generates single-hop question-answer pairs as follows:
Question 1: What is the purpose of nn.LeakyReLU? Answer 1: nn.LeakyReLU is similar to nn.ReLU but allows a small, non-zero slope for negative values, controlled by a "negative_slope" parameter, making it less likely for neurons to become inactive. Question 2: What is the nn.ReLU? Answer 2: nn.ReLU (Rectified Linear Unit) applies the rectified linear unit function element-wise, where any negative input values are set to zero, leaving positive values unchanged.
After that, Multi-hop Question Merging Agent generates the merged question-answering pair as:
Question: How does the behavior of nn.ReLU differ from nn.LeakyReLU in terms of their activation functions when used in a neural network constructed with nn.Sequential? Answer: In a neural network constructed with nn.Sequential, nn.ReLU applies the rectified linear unit function element-wise, where negative input values are set to zero, while positive values remain unchanged.
In contrast, nn.LeakyReLU behaves similarly but allows a small, non-zero slope for negative values through a ’negative_slope’ parameter, which helps prevent neurons from becoming inactive. This means nn.LeakyReLU can potentially maintain more active neurons compared to nn.ReLU in sequential layers.
Further, We try to propose a new prompt for Multi-hop Question Merging Agent to generate more difficult questions for competitive code based on merged question.
Using the following question and answer related knowledge and meanings as a reference, generate additional programming challenges and their solutions specifically designed for high-level coding competitions, such as Olympiad-style problems or advanced AI-related tasks (like those on Kaggle). The focus should be on tasks that require efficient code generation, algorithm design, or AI model implementation.
As shown below, large language models (LLMs) can efficiently produce sophisticated code-generation datasets. This capability represents a compelling and promising research direction. One of generated questions is presented as follows:
Problem Statement: You are tasked with using a genetic algorithm (GA) to optimize the hyperparameters of a neural network for a given task (e.g., image classification using CIFAR-10 dataset). You need to optimize parameters such as the number of layers, the number of neurons in each layer, and the activation function used at each layer. Input: A dataset such as CIFAR-10 or MNIST (for image classification). An initial population of neural network configurations represented by a list of hyperparameters: Number of layers. Number of neurons per layer.
Activation functions (ReLU, LeakyReLU, Sigmoid, etc.).
Further, the corresponding generated answer is shown below:
Output: The optimized neural network configuration after a set number of generations. Solution:
import random
from sklearn.datasets import fetch_openml
from sklearn.model_selection import train_test_split
import torch
import torch.nn as nn
import torch.optim as optim
import torch.nn.functional as F
# Simple feedforward neural network
class NeuralNet(nn.Module):
def __init__(self, layers, activation_fn):
super(NeuralNet, self).__init__()
self.layers = nn.ModuleList()
for i in range(len(layers) - 1):
self.layers.append(nn.Linear(layers[i], layers[i+1]))
self.activation_fn = activation_fn
def forward(self, x):
for layer in self.layers:
x = layer(x)
x = self.activation_fn(x)
return x
# Genetic algorithm for hyperparameter optimization
def genetic_algorithm(dataset, generations=10, population_size=10):
population = []
for _ in range(population_size):
layers = [random.randint(5, 100) for _ in range(random.randint(3, 6))] # Number of neurons in layers
activation_fn = random.choice([nn.ReLU(), nn.LeakyReLU(), nn.Sigmoid()])
population.append((layers, activation_fn))
best_score = float(’inf’)
best_model = None
for generation in range(generations):
for layers, activation_fn in population:
# Convert layers to torch model
model = NeuralNet(layers, activation_fn)
optimizer = optim.Adam(model.parameters())
criterion = nn.CrossEntropyLoss()
# Training loop on dataset
model.train()
optimizer.zero_grad()
# Assume a dummy data loading and training process here
loss = criterion(model(dummy_inputs), dummy_labels)
loss.backward()
optimizer.step()
if loss.item() < best_score:
best_score = loss.item()
best_model = model
# Selection for next generation (randomly choose top models for simplicity)
population = random.sample(population, k=population_size)
return best_model
# Example usage
best_model = genetic_algorithm(fetch_openml("CIFAR_10"))